# MapleGrasp: Mask-guided Feature Pooling for Language-driven Efficient Robotic Grasping

### v1

@@ 1 Introduction
An end-to-end robotic manipulation framework typically comprises four key stages: task planning, visual grounding, object manipulation, and goal-oriented placement. Recent advances have demonstrated the effectiveness of VLMs across these stages, incorporating large-scale pre-trained representations to improve generalization [54]. An alternative paradigm involves Vision-Language-Action (VLA) models, which directly predict trajectories for grasping and manipulating objects, bypassing the modular pipeline [12, 4]. We highlight the difference in these two frameworks in Fig. 1. While VLAs have shown remarkable performance in long-range tasks across both simulation and real-world settings, their deployment in novel environments presents three key challenges: i) Task-Specific Data Collection: Collecting teleoperated or human demonstrations is often difficult and time-consuming for novel interactions [53]; ii) Expensive fine-tuning: Large-scale VLAs require task-specific fine-tuning, which is computationally expensive. Although parameter-efficient fine-tuning (PEFT) techniques alleviate some costs, they restrict adaptability by modifying only a small subset of parameters [20]; iii) Limited generalization: Fine-tuned VLAs often fail to generalize across dynamic environments, particularly when objects are repositioned or when visually similar distractor objects are introduced [46]. Given these challenges, modular frameworks offer an attractive alternative that balances generalizability with ease of deployment. By decoupling perception, planning, and manipulation, modular architectures allow state-of-the-art models in each component to be independently improved and integrated as research advances. Furthermore, LLM-agentic frameworks can serve as high-level decision-makers, dynamically invoking perception and manipulation modules for task execution in novel environments [15, 41, 42]. Compounded errors within modular frameworks can be mitigated using closed-loop feedback and error planning [6, 16].
In this work, we focus on a critical capability within modular frameworks, language-driven grasp detection. This task involves predicting a stable grasp pose for a specified object based on RGB-D images and natural language referring expressions. We focus on 2D grasp detection, which can be naturally extended to 3D using depth information and camera intrinsic parameters. Our contributions include:
@@
    A novel two-stage training strategy for LDRG: initially predicting text-referred object masks, then refining grasp predictions via mask-guided feature pooling.
@@
    Introduce RefGraspNet, a large-scale open-source dataset comprising over 200 million grasp annotations, eight times larger than existing datasets, consisting of challenging real-world scenarios for grasping.
@@
    Extend 2D grasp predictions to 3D via depth maps and inverse kinematics, achieving performance comparable to recent Vision Language Action (VLA) models on the LIBERO benchmark, with superior generalization to diverse tasks while being parameter efficient.
@@ 2 Related Works
Grasp Detection. Robotic grasping with parallel grippers has been extensively studied in both the 2D and 3D domains. The 4-DoF grasp representation defines a grasp using four parameters: $(x,y)$ as the center of the grasp, $w$ as the grasp width, and $\theta$ as the grasp axis angle relative to a top-down hand orientation [1, 8, 17]. The 6-DoF representation extends this to 3D, incorporating three degrees of translation and three degrees of rotation required for stable object grasping [31, 10, 43]. Early approaches to object grasping relied on kinematics and geometric methods to identify stable grasp poses from depth maps [3] and 3D object models, using datasets generated through offline labeled training data, human demonstrations, or trial-and-error methods [13, 9]. The use of RGB-D data has shown promising results for precision grasping, where CNN and ResNet-based architectures leverage visual features to predict grasp pose parameters [22, 48, 65, 58]. Additionally, 3D-CNN models have been employed to extract 3D features from point clouds [39, 38, 51]. Recent advancements in transformer architectures and self-supervised learning have bridged the gap between vision and textual modalities, shifting research towards incorporating language as an additional modality for object manipulation [54]. Language-driven grasp detection (LDGD) refers to the task of associating referring text expressions with RGB-D images to predict a suitable robotic grasp pose for object manipulation [49]. Two primary approaches have emerged for LDGD: i) Visual grounding of referring expressions: This approach predicts segmentation masks for the target object [7, 55, 61], followed by the application of pre-trained grasp detection networks to compute grasp poses [2, 27, 32]. ii) End-to-end grasp pose estimation: This method directly predicts grasp poses from input text and RGB-D modalities [56, 63, 52].
@@
Figure 2: Our framework trains on fused vision-language embeddings for object segmentation, then applies mask pooling in grasp prediction heads. MapleGrasp is an network-agnostic MFP block with no extra parameters and can be used with the CROG backbone as shown.
@@
End-to-End Language-Driven Grasping. Our work focuses on utilizing RGB images of a scene along with free-form referring expressions to determine a stable 4-DoF grasp for the target object in cluttered environments. We extend this by incorporating depth maps to transform the 2D grasp into 3D and leveraging a differential inverse kinematics controller to compute the joint-space trajectory for grasp execution. Several studies have explored similar setups for LDGD [45, 46]. Recent work has introduced diffusion models that learn a probabilistic diffusion process to generate grasp poses during inference [49, 35, 36]. Vision-language fusion provides an alternative approach, offering faster convergence and lower inference times [34, 24, 54]. The ETRG framework enhances vision-language feature fusion for grasp detection by integrating vision-language adapters into frozen CLIP vision and text encoders [57]. MaskGrasp employs a tri-head architecture with cross-attention mechanisms using text embeddings, region-of-interest (ROI) feature embeddings, and mask embeddings [47]. GraspSAM adapts the Segment Anything Model [21] for grasp detection and utilizes GroundingDINO [28] to transform text expressions into object bounding boxes [37].
In our work, we analyze the impact of mask pooling on grasp detection and demonstrate that combining mask pooling with open-set object detectors significantly enhances performance on unseen objects—an issue that previous studies have not adequately addressed. Moreover, we present the first exhaustive comparison of LDRG approaches against large Vision-Language-Action (VLA) models for language-guided grasping tasks.
@@
Vision-Language-Action Models. Direct modeling of vision and language modalities for joint-space robotic control has gained significant interest in recent research. Advances in vision-language models [59, 18, 30], coupled with the availability of large-scale robotic datasets spanning diverse environments, hardware platforms, and tasks [5], have enabled the training of Vision-Language-Action (VLA) models for robotic manipulation [40, 14, 26]. However, the large model sizes of VLAs often demand substantial GPU resources for fine-tuning and real-world deployment [50]. Recent research has sought to mitigate these computational challenges through parameter-efficient training techniques and action chunking methods [19, 20], enabling deployment on smaller GPUs with real-time kernel compatibility for seamless robotic control.
3 Mask-Guided Feature Pooling for Language Driven Grasp Detection
In this section, we introduce our approach of leveraging segmentation masks for feature pooling to focus on relevant regions in an image for targeted grasp pose estimation in cluttered environments. Furthermore, we present our two-stage training framework and discuss our improved training objective used to optimize our method.
@@ 3.1 Language-Driven Grasp Detection
Given an RGB-D image and a corresponding referring expression, language-driven grasping models predict a stable grasp pose to pick up the specified object using either a parallel gripper or a dexterous robotic hand. This task poses three primary challenges: i) accurately identifying the target object using textual cues provided by the referring expression, such as color, shape, and relative positioning; ii) determining a stable and collision-free grasp pose that can be reliably executed in real-world environments without interference from surrounding objects; and iii) generalizing grasp predictions effectively to previously unseen objects that were not included in the training dataset.
We focus on grasp pose estimation in 2D, where where grasps are typically represented as rectangles [49, 37, 34]. Each 2D grasp is defined by the center coordinates $(x,y)$ of the grasp rectangle, the grasp width $w$, and the orientation angle $\theta$ relative to a fixed horizontal axis. Given that this study focuses on top-down grasps performed with a parallel gripper, the height of the grasp rectangle remains constant. Such 2D grasp representations naturally extend into 3D space by incorporating depth information.
Formally, a 2D grasp $\mathbf{g_{2D}}=(x,\,y,\,w,\,\theta)$ is converted into a 3D grasp pose $\mathbf{T_{3D}}$ by projecting the pixel $(x,y)$ into 3D space. With depth value $D(x,y)$ at pixel $(x,y)$, the 3D coordinates relative to the camera are computed as:
$Z=D(x,y),\quad X=\frac{(x-c_{x})\,Z}{f_{x}},\quad Y=\frac{(y-c_{y})\,Z}{f_{y}},$ (1)
where $f_{x},f_{y}$ denote the camera focal lengths, and $c_{x},c_{y}$ represent the principal point offsets. The computed coordinates $(X,Y,Z)$ correspond to the grasp center position in the camera frame. The resulting top-down 3D grasp pose transformation matrix $\mathbf{T}_{3D}$ is then defined as:
$\mathbf{T}_{3D}=\begin{bmatrix}\cos\theta&-\sin\theta&0&X\\ \sin\theta&\cos\theta&0&Y\\ 0&0&1&Z\\ 0&0&0&1\end{bmatrix}.$ (2)
Direct prediction of all four grasp parameters (location, width, and orientation) at every pixel often yields suboptimal results due to their weak direct correlation with the original RGB-D data. Thus, contemporary approaches typically estimate grasp parameters using three distinct 2D prediction maps: i) Grasp Quality Map: Assigns a pixel-wise score for grasp quality; the pixel with the highest score represents the best grasp location, ii) Grasp Width Map: Encodes the optimal grasp width at each pixel and iii) Grasp Angle Map: Stores the grasp orientation $\theta$ at each pixel. This decoupled representation enhances the reliability and interpretability of grasp prediction models.
3.2 Referring Expression Segmentation
Language-Driven Grasp Detection (LDGD) shares strong similarities with a well-established computer vision task known as Referring Expression Segmentation (RES). The goal of RES is to predict a binary segmentation mask that separates the referred object from its surroundings [33, 62]. Since both tasks involve object-level visual grounding, several LDGD frameworks incorporate both segmentation mask prediction and grasp rectangle estimation from the referring query [47, 57]. These architectures typically contain separate projection layers for predicting the binary object mask and the grasp maps. However, since predictions are made independently, the segmentation mask may be correctly identified while grasp maps remain inaccurate.
@@ 3.3 Motivation for Mask Pooling
The predicted binary segmentation masks highlight the image regions corresponding to the referring expression, providing an implicit constraint for grasp pose estimation. If the segmentation mask is correctly predicted, then the optimal grasp pose should lie within the segmented region. This suggests that the grasp prediction module does not need to consider grasp locations outside the segmented object, provided that the segmentation branch can accurately localize the object. With recent advancements in VLMs for RES, zero-shot object segmentation has significantly improved. Consequently, off-the-shelf segmentation models can be used to extract masks for grasp pose estimation networks to predict stable grasp poses [7, 61]. We hypothesize that incorporating segmentation masks as a guiding mechanism within the grasp detection pipeline can improve accuracy by focusing predictions on relevant regions. Prior work has not explored this direction, and we argue that integrating mask pooling with LDRG can effectively utilize advances in RES to enhance grasp detection accuracy.
3.4 Mask-Guided Feature Pooling (MFP)
The concept of utilizing segmentation masks for pixel pooling can be integrated into various LDGD frameworks that jointly predict segmentation masks and grasp maps. We illustrate this integration using the CROG framework [46], a robust LDGD model that leverages CLIP-based visual and textual encoders combined through cross-attention mechanisms to fuse vision-language features. CROG employs a multi-modal Feature Pyramid Network (FPN) followed by additional cross-attention layers that incorporate textual embeddings. These enriched features are subsequently passed through a multi-task projector, which simultaneously outputs binary segmentation masks and grasp parameter maps. The projector architecture consists of sequential Conv2D and upsampling layers designed to maintain feature dimensions consistent with the original input images.
We introduce Mask-Guided Feature Pooling (MFP) within the multi-task projector (see Fig. 2). Our training methodology comprises two sequential stages: i) Pretraining the Segmentation Mask Branch: In this initial stage, we freeze the grasp prediction heads and train the segmentation mask branch separately under a standard Referring Expression Segmentation (RES) training setup. This stage intentionally allows slight overfitting to ensure high-quality object-level masks, facilitating precise pixel pooling in subsequent training. ii) Mask-Guided Grasp Prediction: In this stage, we introduce mask-guided pooling connections within the multi-task projector, focusing the grasp predictions exclusively on visually grounded object regions defined by the segmentation masks. These pooled features undergo further refinement via a simple Multi-Layer Perceptron (MLP) to generate accurate grasp maps.
Two critical design choices ensure robustness. First, we do not use ground-truth object masks for pooling, as this would make grasp predictions overtly sensitive to mask accuracy, which is challenging in cluttered scenes with complex referring text. Second, we avoid using open-set detectors during training, since state-of-the-art object mask segmentation models currently achieve only 50–60% IoU on our benchmarks, which is insufficient for reliable pooling.
3.5 Loss Function
Prior work utilizes the Smooth L1 loss for pixel-wise prediction in visual grounding, penalizing smaller errors quadratically and larger errors linearly, defined per pixel $(\hat{x}_{p},x_{p})$ as:
3.6 Inverse Kinematics Control for Manipulation
Our approach for real world object manipulation transforms the 2D grasp prediction into a desired 3D pose. Specifically, we obtain the final grasp pose in the robot’s base frame, $\mathbf{T}_{\text{base}}^{\text{pose}}$, by composing three transformations:
$\mathbf{T}_{\text{base}}^{\text{pose}}\;=\;\mathbf{T}_{\text{base}}^{\text{ee}}\;\mathbf{T}_{\text{ee}}^{\text{cam}}\;\mathbf{T}_{\text{cam}}^{\text{pose}},$ (4)
where:
  1. 1.
    $\mathbf{T}_{\text{base}}^{\text{ee}}$ is the forward-kinematics transformation from the robot base to its end-effector,
  2. 2.
    $\mathbf{T}_{\text{ee}}^{\text{cam}}$ is the hand–eye calibration matrix (mapping from end-effector to camera),
  3. 3.
    $\mathbf{T}_{\text{cam}}^{\text{pose}}$ is the camera-to-pose transformation predicted by our grasp model.
To execute a grasp in closed loop, we generate a smooth interpolation $\mathcal{C}(t)$ in $\mathrm{SE}(3)$ between the initial end-effector pose $\mathbf{T}_{\text{init}}$ and the desired grasp pose $\mathbf{T}_{\text{goal}}$, as:
$\mathcal{C}(t)\;:\;[0,\,1]\;\to\;\mathrm{SE}(3),\quad\mathcal{C}(0)=\mathbf{T}_{\text{init}},\quad\mathcal{C}(1)=\mathbf{T}_{\text{goal}}.$ (5)
At each control step $t$, the desired pose is
$\mathbf{T}_{\text{des}}\;=\;\mathcal{C}(t).$ (6)
Let $\mathbf{T}_{\text{current}}$ denote the current pose of the end effector. We define the positional error
$\mathbf{e}_{p}=\mathbf{p}_{\text{des}}-\mathbf{p}_{\text{current}},$ (7)
and compute the rotational error using the matrix logarithm of the orientation difference:
$\mathbf{R}_{\text{err}}=\mathbf{R}_{\text{des}}\,\mathbf{R}_{\text{current}}^{\top},\qquad\boldsymbol{\theta}_{\text{err}}=\mathrm{Log}\!\bigl{(}\mathbf{R}_{\text{err}}\bigr{)},$ (8)
where $\mathbf{R}_{\text{des}}$ and $\mathbf{R}_{\text{current}}$ are the rotation components of $\mathbf{T}_{\text{des}}$ and $\mathbf{T}_{\text{current}}$, respectively, and $\mathrm{Log}(\cdot)$ is the matrix logarithm on $\mathrm{SO}(3)$. We form the combined error vector
$\mathbf{e}=\begin{bmatrix}\mathbf{e}_{p}\\ \boldsymbol{\theta}_{\text{err}}\end{bmatrix}\;\in\;\mathbb{R}^{6}.$ (9)
The desired end-effector velocity is then
$\dot{\mathbf{x}}_{\text{target}}\;=\;K_{p}\;\mathbf{e},$ (10)
where $K_{p}$ is a diagonal gain matrix. We compute the joint velocity command by projecting $\dot{\mathbf{x}}_{\text{target}}$ through the robot Jacobian pseudoinverse:
$\dot{\mathbf{q}}_{\text{cmd}}\;=\;\mathbf{J}^{\dagger}\;\dot{\mathbf{x}}_{\text{target}}.$ (11)
Finally, we update the robot’s joint states with $\dot{\mathbf{q}}_{\text{cmd}}$ and advance $t\!\leftarrow\!t+\Delta t$. This process repeats until
$\|\mathbf{T}_{\text{des}}-\mathbf{T}_{\text{current}}\|\;\leq\;\epsilon,$ (12)
indicating successful convergence to the target grasp pose.
4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses

### v2

@@ 1 Introduction
An autonomous robotic manipulation framework typically comprises four key stages: task planning, visual grounding, object manipulation, and goal-oriented placement. Recent advances have demonstrated the effectiveness of VLMs across these stages, incorporating large-scale pre-trained representations to improve generalization [51]. An alternative paradigm involves Vision-Language-Action (VLA) models, which directly predict trajectories for grasping and manipulating objects, bypassing the modular pipeline [11, 3]. We highlight the difference in these two frameworks in Fig. 1. While VLAs have shown remarkable performance in tasks across both simulation and real-world settings, their deployment in novel environments presents three key challenges: i) Task-Specific Data Collection: Collecting teleoperated or human demonstrations is often difficult and time-consuming for novel interactions [50]; ii) Expensive fine-tuning: Large VLAs require task-specific fine-tuning, which is computationally expensive requiring hundreds of GPU hours. Although parameter-efficient fine-tuning (PEFT) techniques alleviate some costs, they restrict adaptability by modifying only a small subset of parameters [18]; iii) Limited generalization: Fine-tuned VLAs often fail to generalize in dynamic environments, particularly when objects are repositioned or when visually similar distractor objects are introduced [44]. Given these challenges, modular frameworks offer an attractive alternative that balances generalizability with ease of deployment. By decoupling perception, planning, and manipulation, modular architectures allow state-of-the-art models in each component to be independently improved and integrated as research advances. Furthermore, LLM-agentic frameworks can serve as high-level decision-makers, dynamically invoked perception and manipulation modules for task execution in novel environments [13, 36, 38]. Compounded errors within modular frameworks can be mitigated using closed-loop feedback and error planning [5, 14].
In this work, we focus on a critical capability within modular frameworks, language-driven grasp detection. This task predicts a stable grasp pose for a specified object based on RGB-D images and a free-form language expression. An accurate LDRG model can enable low-level execution for high-level LLM planners, advancing fully autonomous agentic frameworks for human-robot interaction. Our proposed method, MapleGrasp, can predict both 4 DoF top-down and oriented 6 DoF grasps. Our contributions include:
@@
    A novel two-stage training architecture for LDRG: initial text-referred object mask prediction, followed by mask-guided feature pooling for grasp refinement.
@@
    Introduce RefGraspNet, a large-scale open-source benchmark comprising over 200 million grasps, eight times larger than existing datasets (Tab. 1), consisting of challenging real-world scenarios for grasping.
@@
[…]
@@ 2 Related Works
Grasp Detection. Robotic grasping with parallel grippers has been extensively studied in both the 2D and 3D domains. The 4-DoF grasp representation defines a grasp using four parameters: $(x,y)$ as the center of the grasp, $w$ as the grasp width, and $\theta$ as the grasp axis angle relative to a top-down hand orientation [1, 7, 15]. The 6-DoF representation extends this to 3D, incorporating three degrees of translation and three degrees of rotation required for stable object grasping [29, 9, 39]. The use of RGB-D data has shown promising results for precision grasping, where CNN and ResNet-based architectures leverage visual features to predict grasp pose parameters [20, 46, 63, 56]. Recent advancements in transformer architectures and self-supervised learning have bridged the gap between vision and textual modalities, shifting research towards incorporating language as an additional modality for object manipulation [51]. Language-driven grasp detection (LDGD) refers to the task of associating referring text expressions with RGB-D images to predict a suitable robotic grasp pose for object manipulation [47]. Two primary approaches have emerged for LDGD: i) Visual grounding of referring expressions: Predicts segmentation masks for the target object [6, 53, 60], followed by pre-trained grasp detection networks to compute grasp poses [2, 25, 30]. ii) End-to-end grasp pose estimation: Directly predicts grasp poses from input text and RGB-D images [54, 61, 49].
@@
Figure 2: MapleGrasp: Our framework trains on fused vision-language embeddings for object segmentation, then applies mask pooling in grasp prediction heads. The predicted grasp maps can be used for both top-down and oriented grasping.
@@
End-to-End Language-Driven Grasping. Recent work has introduced diffusion models that learn a probabilistic diffusion process to generate grasp poses during inference [47, 32, 33]. Vision-language fusion provides an alternative approach, offering faster convergence and lower inference times [31, 22, 51]. The ETRG framework enhances vision-language feature fusion for grasp detection by integrating trainable adapters into frozen CLIP vision and text encoders [55]. MaskGrasp employs a tri-head architecture with cross-attention mechanisms using text embeddings, region-of-interest feature embeddings, and mask embeddings [45]. GraspSAM adapts the Segment Anything Model [19] for grasp detection and utilizes GroundingDINO [26] to transform text expressions into object bounding boxes [34]. GraspMolmo [8] trains the Molmo VLM on the grasp point detection task followed by a grasp proposal sampler to identify accurate grasps for task oriented grasping. Our proposed architecture predicts grasp pose using quality, width and angle maps. Several studies have explored similar setups for LDGD [43, 44, 41]. These maps can be used to predict both 4 DoF and 6 DoF grasps by computing the grasp point on the object and overlapping with an independent grasp proposal network.
In this work, we study mask pooling for grasp detection and demonstrate that localizing the grasp prior to predicting its angle and width leads to a more efficient learning process. This approach improves both training efficiency and test-time accuracy. Additionally, we present the first comprehensive comparison of LDRG approaches with large VLAs in the context of language-guided grasping tasks.
@@
Vision-Language-Action Models. Direct modeling of vision and language modalities for joint-space robotic control has gained significant interest in recent research. Advances in vision-language models [58, 16, 28], coupled with the availability of large-scale robotic datasets spanning diverse environments, hardware platforms, and tasks [4], have enabled the training of Vision-Language-Action (VLA) models for robotic manipulation [35, 12, 24]. However, the large model sizes of VLAs often demand substantial GPU resources for fine-tuning and real-world deployment [48]. Recent research has sought to mitigate these computational challenges through parameter-efficient training techniques and action chunking [17, 18], enabling deployment on smaller GPUs with real-time kernel compatibility.
3 Methodology
We propose object-mask feature pooling to isolate relevant regions for grasp detection in clutter, detail our architecture (Fig. 2), and outline an enhanced training objective.
@@ 3.1 Language-Driven Grasp Detection
From an RGB-D image and a referring expression identifying a target object to grasp, language-driven grasping models must predict a stable grasp pose for a parallel gripper or dexterous hand. This poses three primary challenges: (i) identifying the target from textual cues (color, shape, spatial relations); (ii) selecting a collision-free grasp executable in real settings; and (iii) generalizing to unseen objects. Vision-language driven grasping is studied with both 4 DoF top-down and 6 DoF oriented grasps. While 6 DoF grasps are more accurate, annotating such grasp poses at a large scale to create real-world datasets is challenging. A number of works thus rely on image-based datasets with 4 DoF grasp rectangles. We design our method such that it can be used with both 4 DoF and 6 DoF grasping datasets. Fig. 3 illustrates the two grasp formats and their transformation via an external proposal sampler.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Conversion from a 4 DoF to a 6 DoF grasp pose: (top-left) initial top-down grasp rectangle, (top-right) sampled contact points reprojected to 2D, (bottom-left) 6 DoF grasp proposals from Contact-GraspNet [40], (bottom-right) selected 6 DoF grasp.
Formally, MapleGrasp predicts three dense maps, each of size $H\times W$: a grasp quality map $Q\in\mathbb{R}^{H\times W}$, a grasp width map $w\in\mathbb{R}^{H\times W}$, and a grasp angle map $\theta\in\mathbb{R}^{H\times W}$. The quality map assigns a score to each pixel, indicating the likelihood of a successful grasp. The width map encodes the optimal gripper width at each pixel, while the angle map represents the in-plane grasp orientation $\theta\in[-\pi/2,\pi/2]$. A 2D rectangular grasp $g_{\text{2D}}=(x^{*},y^{*},w^{*},\theta^{*})$ is extracted by selecting the pixel $(x^{*},y^{*})=\arg\max_{(i,j)}Q_{i,j}$, and reading $w^{*}=w_{x^{*},y^{*}}$, $\theta^{*}=\theta_{x^{*},y^{*}}$. This rectangular grasp is projected onto the point cloud $\mathcal{P}$ to identify a subset of points $\mathcal{P}_{g}\subset\mathcal{P}$ within the grasp rectangle. A set of candidate 6-DoF grasp poses $\{G_{k}\}$ is generated using an independent grasp sampler. Each candidate pose $G_{k}$ is evaluated by computing its intersection with $\mathcal{P}_{g}$, and the final grasp $G^{*}$ is chosen as the one that overlaps the maximum number of graspable points: $G^{*}=\arg\max_{k}\,|\mathcal{P}_{g}\cap\mathcal{V}(G_{k})|$, where $\mathcal{V}(G_{k})$ denotes volume swept by the gripper pose $G_{k}$.
@@ 3.2 Motivation for Mask Pooling
Most LDRG methods predict an object mask and grasp pose simultaneously. A binary object segmentation mask already pinpoints the pixels referenced by the query, implicitly limiting where a valid grasp can lie. If the mask is correct, the grasp network need not evaluate locations outside it, shrinking search space and reducing errors. We treat the mask as an explicit guiding mechanism: features are pooled within the segmented region before grasp prediction.
3.3 MapleGrasp
Given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a referring instruction $T$, we compute vision-language embeddings using a pre-trained CLIP model: $x_{\text{img}}=\text{CLIP}_{\text{img}}(I)$, $x_{\text{text}}=\text{CLIP}_{\text{text}}(T)$. The visual embedding $x_{\text{img}}$ is passed through a multi-modal Feature Pyramid Network $\mathcal{F}$, and fused with the text embedding via cross-attention, resulting in $x_{\text{fused}}=\text{CrossAttn}(\mathcal{F}(x_{\text{img}},x_{\text{text}}),x_{\text{text}})$. These embeddings are then processed through a series of convolutional and upsampling layers to produce a spatial feature map $f\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$, where $H^{\prime}$ and $W^{\prime}$ are intermediate spatial resolutions and $C$ is the channel dimension. We introduce Mask-Guided Feature Pooling within a two-stage training framework to guide grasp prediction. In Stage I (Segmentation Training), we freeze the grasp prediction heads and train the segmentation branch independently. The text embedding $x_{\text{text}}$ is projected via a linear layer to obtain $z\in\mathbb{R}^{C}$, which is then used to compute the object-centric mask via a per-pixel dot product: $M_{i,j}=\sigma(f_{i,j}^{\top}z)$, yielding a soft binary mask $M\in[0,1]^{H^{\prime}\times W^{\prime}}$. This stage allows mild overfitting to ensure accurate object localization for downstream pooling. In Stage II (Mask-Guided Grasp Prediction), we train the full model end-to-end. The predicted mask $M$ is used to modulate the visual features via element-wise multiplication, producing $f_{\text{pooled}}=f\odot M$. These pooled features are refined and passed through three parallel projection heads to predict the grasp quality $Q=\mathcal{H}_{q}(f_{\text{pooled}})$, angle $\theta=\mathcal{H}_{\theta}(f_{\text{pooled}})$, and width $w=\mathcal{H}_{w}(f_{\text{pooled}})$.
Our goal was to support both 4-DoF and 6-DoF grasping with minimal architectural changes for future dataset compatibility. For image-only datasets with 4-DoF grasp pose annotations, we utilize the grasp rectangle to compute ground truth quality, angle, and width maps, following prior work [44]. For RGB-D datasets with 6-DoF grasp poses, we first identify the 3D area on the target object where ground truth poses occur. This region is then projected into 2D to generate a rectangle over the graspable area of the object, which is subsequently treated in a manner analogous to the 4-DoF case. MapleGrasp computes the graspable region on the object, which is then overlapped with candidate 6-DoF grasp poses generated by an off-the-shelf grasp sampler, Contact-GraspNet [40]. The candidate with maximum overlap is selected as the collision-free grasp for the target.